OpenStudy (anonymous):

What is 4.6 × 108 cm/min in meters per hour? A. 2.76 × 108 m/h B. 2.76 × 1010 m/h C. 7.67 × 106 m/h D. 7.67 × 108 m/h

OpenStudy (anonymous):

@EngrChong

OpenStudy (anonymous):

298.08 meters per hour

OpenStudy (anonymous):

4.6*108 =?

OpenStudy (anonymous):

100cm =1 meter

OpenStudy (anonymous):

i still don't understand what the answer gonna be?

OpenStudy (anonymous):

the answer is not in your options

OpenStudy (anonymous):

has to be

OpenStudy (anonymous):

its 298.08 meters per hour

OpenStudy (anonymous):

WTF this is for school and it's online how the hell is the answers i posted are wrong?

OpenStudy (anonymous):

how do convert min to hour and cm to meter

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

1 hour =60 min 1min=1/60 then 1 cm=0.001 meter 1meter =100cm

OpenStudy (anonymous):

1cm =1/100

OpenStudy (anonymous):

but i need my A,B,C or D answer ASAP

OpenStudy (anonymous):

ok i will explain .

OpenStudy (anonymous):

ughhhh no need to explain i just need the answer

OpenStudy (anonymous):

4.6*108/100*60 so 4.6*18/100*10 8.28/1000 so

OpenStudy (anonymous):

NOT HELPING

OpenStudy (anonymous):

see 8.28 can written as

OpenStudy (anonymous):

its against rules saying the answer @wormie2014 .u cant get it if i say u the answer

OpenStudy (anonymous):

IDGAF i just need it...

OpenStudy (anonymous):

u tell me 8.28/1000 can written as.here is ur answer

OpenStudy (anonymous):

that slash means divide right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0.00828 is the answer i got but it's not any of the choices.

OpenStudy (anonymous):

wtf and idgaf what do u mean by that @wormie2014 .dont use these kinds of words here.you are so rude to get ur answer done

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

i think the question has a typo, it should be : 4.6 × 10^8 cm/min

OpenStudy (anonymous):

no i got the answer @ganeshie8

OpenStudy (anonymous):

@wormie2014 needs only the answer.

OpenStudy (anonymous):

i'm sorry for being so rude @dg2 im just really aggervated right now

OpenStudy (trinibutera):

SO WHATS THE ANSWER?